The sales of a company from 1993-1998 are given below:

Year 1993 1994 1995 1996 1997 1998
Sales (in lakhs of rupees) 40 45 50 55 60 65
Fit a linear curve using the least squares method. Hence find out the company’s sales in 1999.
1
Expert's answer
2020-06-11T19:15:56-0400

Let "x=" the number of years since 1993, "y=" sales (in lakhs of rupees)


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & x & y & xy & x^2 & y^2 \\ \hline
 & 0 & 40 & 0 & 0 & 1600 \\
 \hdashline
 & 1 & 45 & 45 & 1 & 2025 \\
 & 2 & 50 & 100 & 4 & 2500\\
 \hdashline
 & 3 & 55 & 165 & 9 & 3025\\
 & 4 & 60 & 240 & 16 & 3600 \\
 & 5 & 65 & 325 & 25 & 4225\\
 \hdashline
 
 Sum= & 15 & 315 & 875 & 55 & 16975
\end{array}"

"\bar{x}={1\over n}\displaystyle\sum_{i=1}^nx_i=2.5, \bar{y}={1\over n}\displaystyle\sum_{i=1}^ny_i=52.5"

"S_{xx}=\displaystyle\sum_{i=1}^nx_i^2 -{1\over n}(\displaystyle\sum_{i=1}^nx_i)^2=17.5"

"S_{yy}=\displaystyle\sum_{i=1}^ny_i^2 -{1\over n}(\displaystyle\sum_{i=1}^ny_i)^2=437.5"

"S_{xy}=\displaystyle\sum_{i=1}^nx_iy_i -{1\over n}(\displaystyle\sum_{i=1}^nx_i)(\displaystyle\sum_{i=1}^ny_i)=87.5"

"m={S_{xy}\over S_{xx}}={87.5\over 17.5}=5"

"n=\bar{y}-m\cdot\bar{x}=52.5-5\times2.5=40"

Therefore, we find that the regression equation is:


"y=40+5x"




Find out the company’s sales in 1999


"y=40+5\times6=70(lakhs\ of\ rupees)"

The company’s sales in 1999 will be 70 lakhs of rupees.